The average monthly salary for first here teachers is 21,945 if the distribution is approximately normal with a standard deviation of 3250 how much will a teacher earn in a month if his salary has a z score of 1.15


1
Expert's answer
2021-05-11T07:03:49-0400

We have given that,

"\mu = 21945"

"\sigma = 3250"

"z = 1.15"

We know that,

"z = \dfrac{x-\mu}{\sigma}"


"1.15 = \dfrac{x-21945}{3250}"


"1.15 \times 3250 = x- 21945"


"x = 25682.5"